OpenStudy (anonymous):

What are the solutions, if any, of the following equation? 6y − 3 = −6y + 3 A. There is no solution. <--- B. There are an infinite number of solutions. C. 0.5 D. 2 @hartnn

jimthompson5910 (jim_thompson5910):

That is not correct.

hartnn (hartnn):

^ add 6y on both sides first

OpenStudy (pinksapphire):

If it is 0=0 then there are infinite amount of solutions but if there is like 6=3 then there are no solutions.

OpenStudy (pinksapphire):

Do you get it? The answer will be B...

OpenStudy (anonymous):

Sorry! It was a typo. I asked my sister to type it in. She put A not B like I asked. :)

OpenStudy (anonymous):

If you look at it, both of the equations are the same. So there will be infinite solutions

OpenStudy (loser66):

@jim_thompson5910 Please, make it clear, why can't I get the answer like others? They said it is infinite solutions while my solution is 0.5.This is my work 6y -3 =-6y +3 +6y both sides 6y -3+6y =3 12y -3 =3 +3 both sides 12y-3+3 = 3+3 12y =6 divide both sides by 12 y = 6/12 = 0.5 Please , check what is wrong with my solution.

jimthompson5910 (jim_thompson5910):

y = 0.5 or y = 1/2 is correct

jimthompson5910 (jim_thompson5910):

it is the only solution

OpenStudy (loser66):

oh, thank you :)

jimthompson5910 (jim_thompson5910):

np